OpenStudy (anonymous):

VECTORS: HELP ME A and B have [position vectors \(3i+2j-k\) and \(2i-j-8k\) respectively. a. Find: i. AB ii. the unit vector \(u\) in the direction of BA. b. Is \(u\) perpendicular to OA? c. If C has osition vector \(i+j+ak\), and OC is perpendicular to \(ai-j+4k\), find \(a\). d. If M is the midpoint of [AB], find the position vector of M. e. Line

OpenStudy (anonymous):

vector AB=(2i-j-8k)-(3i+2j-k)=-i-3j-7k \[AB=\sqrt{\left( -1 \right)^2+\left( -3 \right)^2+\left( -7 \right)^2}=\sqrt{59}\]

OpenStudy (anonymous):

i have to go now.

OpenStudy (anonymous):

shouldn't AB be B-A?

OpenStudy (anonymous):

i mean the one you did was the magnitude, correct?

OpenStudy (anonymous):

vector AB=position vector of B-position vector of A

OpenStudy (anonymous):

vector BA=i+3j+7k unit vector u in the direction of BA\[=\frac{ i+3j+7k }{ \sqrt{1^2+3^2+7^2} }=\frac{ i+3j+7k }{ \sqrt{59} }\]

OpenStudy (anonymous):

\[OA.AB=\left( 3i+2j-k \right).\left( -i-3j-7k \right)=3(-1)+2(-3)+(-1)(-7)\] \[=-3-6+7=-2\pm \neq 0\] OA is not perpendicular to AB

OpenStudy (anonymous):

C \[\left( i+j+ak \right).\left( ai-j+4k \right)=0\] 1(a)+1(-1)+a(4)=0 5a=1 a=1/5

OpenStudy (anonymous):

D P.V. of M \[=\frac{ (3i+2j-k)+(2i-j-8k) }{ 2 }=\frac{ 5i+j-9k }{ 2 }\]

OpenStudy (anonymous):

E. eq. of line AB is \[vector~ r=3i+2j-k +\lambda \left( -i-3j-7k \right)\]